0 0
Add a comment Improve this question Transcribed image text
Answer #1

(a)

volume at T = 26 K,

moles of helium,

n = 10.1 / 4 = 2.525 moles

according to gas equation,

P*V = n*R*T

V1 = n*R*T1 / P

V1 = 2.525*8.314*26 / 1.013*10^5

V1 = 5.38*10^(-3) m^3

(b)

volume at T = 295 K

V2 = n*R*T2 / P

V2 = 2.525*8.314*295 / 1.013*10^5

V2 = 6.11*10^(-2) m^3